OpenStudy (anonymous):

What are the x– and y–intercepts of the graph of y = x2 – 10x + 21?

jhonyy9 (jhonyy9):

x1,2=(10+/-sqrt(100-84))/2=(10+/-4)/2=6/2 =3 and 14/2 =7 so than x=3 and x=7 are these zeros for function y